OpenStudy (anonymous):

@johnweldon1993 @campbell_st The diameter of a circle is 22 inches. What is the radius of the circle? A. 11 in. B. 33 in. C. 44 in. D. 66 in.

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

HELP please

OpenStudy (campbell_st):

well a radius is half the diameter... hope it helps

hero (hero):

\(r = \dfrac{d}{2}\)

OpenStudy (anonymous):

So 66pi

OpenStudy (_michaeld22):

@campbell_st and @Hero described it perfectly. The radius is half of the diameter.

OpenStudy (anonymous):

so divide 66 and 2

OpenStudy (campbell_st):

the diameter given in the question is 22 in what is half of that measure...?

OpenStudy (_michaeld22):

You said that the diameter is 22 inches, right? So the radius is 22/2.

OpenStudy (anonymous):

that equal 22/2=11

OpenStudy (_michaeld22):

Yup, which is answer A.

OpenStudy (anonymous):

Yay what did i win LOL

OpenStudy (_michaeld22):

XD Hmm. A correct answer?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

can u guys help with more

OpenStudy (_michaeld22):

Ehh I'm not very good at this kind of stuff. But yeah, sure, I'll try to help.

OpenStudy (anonymous):

Great

OpenStudy (anonymous):

A circular patio has a radius of 18 feet. What is the longest distance across the patio? A. 9 ft B. 18 ft C. 36 ft D. 54 ft

OpenStudy (anonymous):

18/2=9

OpenStudy (_michaeld22):

So you basically reverse the other question. \[d = 2r\]

OpenStudy (anonymous):

am i right?

OpenStudy (_michaeld22):

So 2*18 = ?

OpenStudy (_michaeld22):

You gotta multiply this one, not divide.